# Chinese Academy of Sciences, Beijing 100190, China

### v1

@@ 5.1.2 Effects on the neutrino Boltzmann hierarchies
In order to obtain the accurate CMB spectral power spectrum and the matter power spectrum, we need to solve a coupled set of the Boltzmann-Einstein equations for the perturbations of the metric, photon, baryon, cold dark matter and massive neutrinos. Here we briefly review how neutrino asymmetries and spectral distortions affect the Boltzmann-Einstein equations. Following Ref. Ma:1995ey , the perturbed distribution functions for the three mass eigenstate neutrinos can be written as
@@
$ds^{2}=a^{2}(\tau)\{-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\}.$ (91)
The perturbation of the metric $h_{ij}$ can be decomposed as $h_{ij}=h\delta_{ij}+h_{ij}^{\parallel}+h_{ij}^{\perp}+h_{ij}^{T}$, where $h\equiv h_{ii}$ is the trace part of $h_{ij}$ , and $h_{ij}^{\parallel}$, $h_{ij}^{\perp}$ and $h_{ij}^{T}$ are the traceless parts, satisfying
@@
By definition, $h_{ij}^{\parallel}$ and $h_{ij}^{\perp}$ can be written in terms of a scalar filed $\mu$ and a divergenceless vector $\vec{A}$,
$h_{ij}^{\parallel}$ $=\left(\partial_{i}\partial_{j}-\frac{1}{3}\delta_{ij}\nabla^{2}\right)\mu$ (93)
@@
Therefore, the scalar modes of the metric perturbations are characterised by the two scalar fields $h$ and $\mu$, while the vector and the tensor modes are characterised by $\vec{A}$ and $h_{ij}^{T}$, respectively.
In this work we will only consider the scalar modes of $h_{ij}$, which can be Fourier transformed as
@@
where $h(\vec{k},\tau)$ and $\eta(\vec{k},\tau)$ are the trace part and the traceless part of the scalar modes in the Fourier space which are related to the scalar fields $h$ and $\mu$ in the real space. Since neutrinos are collisionless after decoupling, the evolution of $\Psi_{i}$ obeys the first-order collsionless Boltzmann equation in the Fourier-space
@@
where $\epsilon\equiv aE=\sqrt{q^{2}+a^{2}m^{2}_{\nu_{i}}}$ is the comoving energy. Since Eq. (95) is independent of the azimuth angle, we can expand $\Psi$ in a Legendre series
@@
Inserting Eq. (96) into Eq. (95) and using the orthonormality of the Legendre polynomials, we can obtain the Boltzmann hierarchies for $\Psi_{i,l}$:
@@
Note that the effect of neutrino asymmetries and spectral distortions is manifested by the neutrino density distribution term $\frac{{d}\ln\varrho_{i}^{0}}{{d}\ln q}$. Since the density matrices for neutrinos and antineutrinos with different flavors are different in the presence of primordial neutrino asymmetries, the Boltzmann hierarchies for the corresponding perturbations $\Psi_{i}$ and $\bar{\Psi}_{i}$ are in general different and should be treated separately. With the moments $\Psi_{i,l}$, the perturbations in physical quantities, such as the energy density, pressure, energy flux and shear stress, can be computed straightforwardly by integrating different moments. For our purpose, the fractional density perturbation for the neutrinos with family number $i$ can be expressed as
@@
where $\rho_{i,0}(t)$ is the zeroth-order homogeneous background energy density for neutrinos. Eventually, the effects of neutrino masses and asymmetries at different length scales can be reflected in the evolution of the Fourier modes $\delta_{i,k}$, which we obtain via the CLASS code.
@@ 5.1.3 Effects on the BAO
Following Ma:1995ey , the evolution of the perturbations for photon obey the Boltzmann hierarchy,
$\dot{\delta}_{\gamma}$ $=$ $-{4\over 3}\theta_{\gamma}-{2\over 3}\dot{h}\,,$
@@
$\dot{F}_{\gamma\,l}$ $=$ ${k\over 2l+1}\left[lF_{\gamma\,(l-1)}-(l+1)F_{\gamma\,(l+1)}\right]-an_{e}\sigma_{T}F_{\gamma\,l}\,,\quad l\geq 3\,,$
@@
where $\delta_{\gamma}$ is the density fluctuation of photon, $\theta_{\gamma}$ is the divergence of photon velocity, and ${F}_{\gamma\,l}$ is the density perturbation of photon. In addition, $n_{e}$ is the electron number density and $\sigma_{T}$ is the Thomson scattering cross section. The Boltzmann equations for baryon are
$\dot{\delta}_{b}$ $=$ $-\theta_{b}-{1\over 2}\dot{h}\,,$
$\dot{\theta}_{b}$ $=$ $-{\dot{a}\over a}\theta_{b}+c_{s}^{2}k^{2}\delta_{b}+{4\bar{\rho}_{\gamma}\over 3\bar{\rho}_{b}}an_{e}\sigma_{T}(\theta_{\gamma}-\theta_{b})\,,$ (100)
where $\delta_{b}$ and $\theta_{b}$ is the baryon density fluctuation and the divergence of baryon fluid velocity, and $c_{s}^{2}$ is the square of the baryon sound speed.
Due to the electromagnetic interactions between photons and electrons, they evolve differently before and after recombination. Before recombination, photons and baryons are tightly coupled so that they can be treated as a single photon-baryon fluid, called as the "tight-coupling approximation". During such "tight-coupling epoch", the perturbations of baryon and photon undergo harmonic motion with very slowly decaying amplitude (Peebles:1970ag, ; Sunyaev:1970bma, ). After recombination, photons release their hold on the baryons, so that the baryon acoustic oscillations (BAO) become frozen-in, and the the baryon perturbations begin to grow as $\propto a$. The BAO feature then gets transferred to the gravitational potential and the matter fluctuations, leaving imprints in the galaxy power spectrum that depends on the sound horizon at decoupling and the angular diameter distance to the surface of photon-electron last-scattering.
In the presence of primordial neutrino asymmetries, the matter-radiation equality will be directly modified by the additional contribution to $N_{\rm eff}$, leading to the modification of the BAO properties. In addition, the effects of primordial neutrino asymmetries on the evolution of neutrino perturbations are also transported to photons and baryons through metric perturbations encoded in $h$ and $\eta$, even though neutrinos are collisionless. As a consequence, the primordial neutrino asymmetries have significant implications for both the CMB angular power spectrum and the BAO in the matter power spectrum, as we will see later.
@@ 5.2 Results of the CMB anisotropies and the matter power spectrum
Figure 7: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ within the SM case for different total neutrino masses, compared to the results with massless neutrinos. Right panel: The linear matter power spectrum $P(k)$ within the SM case for different total neutrino masses, compared to the results with massless neutrinos.
@@
We present in this subsection the numerical results of the CMB and LSS obtained via the CLASS code. Since both observable are affected by the neutrino masses and asymmetries at the same time, we first discuss the effect of neutrino masses on the CMB and LSS within the SM. We present the ratio of the CMB angular power spectrum between the results with massive neutrinos and massless neutrinos $C_{l}^{\rm TT}/C_{l}^{\rm TT}(\sum m_{\nu}=0)$ and the similar ratio for the matter power spectrum $P_{k}/P_{k}(\sum m_{\nu}=0)$ for several choices of total neutrino masses in the left and right panels of Fig. 7, respectively. Note that the normal order of the neutrino mass hierarchy is considered in this work, so we adjust the $\mu_{\nu_{1}}$ for each choice of total mass, while $\mu_{\nu_{2}}$ and $\mu_{\nu_{3}}$ are obtained by the fixed mass differences between neutrinos. Focusing on the left panel, the ratio of the $C_{l}^{\rm TT}$ between massive and massless neutrinos can be divided into three stages. At large angular scales (lower multipoles ${\ell\lesssim 20}$), massive neutrinos induce a small suppression which comes from the late time integrated Sachs-Wolfe effect. On the intermediate scales (${20\lesssim\ell\lesssim 500}$), dominated by the imprints of the acoustic oscillations in the baryon-photon fluid, the first peak of $C_{l}^{\rm TT}$ is affected by the modified evolution of the gravitational potential due to the correlation between massive neutrinos and metric perturbations. As we can see, the height of the first peak given by the early integrated Sachs-Wolfe effect is increased for larger neutrino masses compared to the case with massless neutrinos. The acoustic peaks at small scales (${\ell\gtrsim 500}$) are also affected by the massive neutrinos. Similar with the first peak, the height of the acoustic peaks is increased for larger neutrno masses.
@@
Similarly, the resulting matter power spectrum with massive neutrinos can also be divided to three stages. At large scales corresponding to $k\lesssim 10^{-3}\,h/{\rm Mpc}$, neutrinos do not enter the horizon, and the small suppression of $P(k)$ is induced by the modification of the matter-radiation equality from the non-relativistic transition of massive neutrinos. At intermediate scales corresponding to $10^{-3}\,h/{\rm Mpc}\lesssim k\lesssim 1\,h/{\rm Mpc}$, neutrinos gradually enter the horizon, so that the free streaming of these neutrinos results in an additional suppression of the matter power spectrum. Therefore, $P_{k}/P_{k}(\sum m_{\nu}=0)$ decreases with increasing wavenumber $k$. And at small scales ($k\gtrsim 1\,h/{\rm Mpc}$), the majority of neutrinos stream freely, leading to an almost constant suppression of the power spectrum. The larger the total neutrino mass, the greater the suppression of the matter power spectrum. It is also worth noting that the matter power spectrum is more sensitive to the mass of the neutrinos, as can be seen from the plots. Therefore, it is more efficient to constrain the neutrino masses with the LSS observations, as we will see in the next section.
@@
Figure 8: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the results with $\xi_{\nu}=0$. Right panel: The linear matter power spectrum $P(k)$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the results with $\xi_{\nu}=0$. The total neutrino mass is fixed to $\sum m_{\nu}=0.1{\rm eV}$ for both panels.
@@
Figure 9: The sound horizon at the baryon drag epoch of the early universe $r_{d}$ as a function of the primordial neutrino asymmetry $\xi_{\nu}$.
@@
We now present the results for non-zero $\xi_{\nu}$. To isolate the effects of neutrino asymmetries from the total neutrino mass, the left panel of Fig. 8 shows the ratio of the CMB angular power spectrum for $\xi_{\nu}\neq 0$ to $\xi_{\nu}=0$, with $\sum m_{\nu_{i}}=0.1\,\text{eV}$. A similar ratio for the matter power spectrum is shown in the right panel. As seen in the left panel, the effects of neutrino asymmetries on the CMB angular power spectrum manifest primarily at small scales ($\ell\gtrsim 400$). At $\ell\lesssim 400$, $C_{l}^{\rm TT}$ for non-zero $\xi_{\nu}$ matches the results for $\xi_{\nu}=0$. However, as $\ell$ increases, deviations due to $\xi_{\nu}$ become apparent. These deviations depend on both the magnitude and sign of $\xi_{\nu}$: positive $\xi_{\nu}$ enhances the small-scale power, while negative $\xi_{\nu}$ leads to suppression. This behavior is largely driven by changes in helium abundance, which alters the CMB spectrum via diffusion damping. In addition to overall enhancement or suppression, the acoustic peaks are also affected by $\xi_{\nu}$, as larger $\xi_{\nu}$ values modify the oscillations at the last scattering surface. The amplitude of these modifications grows with the absolute value of $\xi_{\nu}$, with negative $\xi_{\nu}$ producing more pronounced effects.
@@
For the linear matter power spectrum in the right panel of Fig. 8, we observe that at large scales, the result for non-zero $\xi_{\nu}$ is identical to that for $\xi_{\nu}=0$. At small scales, however, the free-streaming of neutrinos is enhanced with larger $\xi_{\nu}$, resulting in a slightly stronger suppression of the matter power spectrum, although this suppression is insensitive to the sign of $\xi_{\nu}$. The more intriguing effect of non-zero $\xi_{\nu}$ is the oscillations of the matter power spectrum on the scales $10^{-1}\,h/{\rm Mpc}\gtrsim k\gtrsim 10^{0}\,h/{\rm Mpc}$, which represents the effect of the BAO. Similar to the acoustic oscillations in the CMB angular power spectrum, we find that the amplitudes of the BAO in the matter power spectrum are also larger with a larger absolute value of $\xi_{\nu}$, while the negative $\xi_{\nu}$ leads to a larger effect than that for the positive $\xi_{\nu}$. In addition to the imprints of BAO on the matter power spectrum, we show in Fig. 9 the sound horizon at the baryon drag epoch of the early universe $r_{d}$ as a function of the primordial neutrino asymmetry $\xi_{\nu}$. Due to the additional contribution to $N_{\rm eff}$ for non-zero $\xi_{\nu}$ according to Eq. 63, the larger the $\left|\xi_{\nu}\right|$, the smaller the $r_{d}$, resulting in the significant imprints on the matter power spectrum. These results imply that BAO experiments can also provide powerful constraints on the primordial neutrino asymmetry, as we will in Section 6, which is often overlooked in previous studies. Note also that these conclusions are drawn with a fixed total neutrino mass $\sum m_{\nu_{i}}=0.1\,\text{eV}$. However, they remain valid for different total neutrino masses, although the resulting ratios $C_{l}^{\rm TT}/C_{l}^{\rm TT,\,\xi_{\nu}=0}$ and $P(k)/P^{\xi_{\nu}=0}(k)$ may vary slightly at intermediate scales due to the non-relativistic transition of massive neutrinos.
@@
Figure 10: The CMB angular power spectrum $C_{l}^{\rm TT}$ (left panel) and the linear matter power spectrum $P(k)$ (right panel) for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the corresponding results with $\xi_{\nu}=0$. The total neutrino mass is fixed to $\sum m_{\nu}=0.1{\rm eV}$ for both panels. The solid curves are the same with Fig. 5, corresponding to the results including the spectral distortions of the neutrinos and antineutrinos, while the dashed curves correspond to the results without the spectral distortions.
@@
Finally, we briefly discuss the effects of spectral distortions of the neutrinos and antineutrinos on the CMB and LSS. Similar to Fig. 8, we present the resulting CMB angular power spectrum $C_{l}^{\rm TT}$ and linear matter power spectrum $P(k)$ with non-zero $\xi_{\nu}$, neglecting the neutrino spectral distortions, as the dashed curves in the left and right panels of Fig. 10. The solid curves in Fig. 10 are the same as those in Fig. 8, corresponding to the results including the spectral distortions of the neutrinos and antineutrinos. Comparing the resulting CMB angular power spectrum $C_{l}^{\rm TT}$ with and without the neutrino spectral distortions, we see that the primary effect of including the spectral distortions is a slight suppression of the first three acoustic peaks, with an almost negligible effect on the remaining peaks in the small scales. For the linear power spectrum, we find that the effect of spectral distortions on the BAO is similar with that on the CMB angular power spectrum. Here the first two acoustic peaks are slightly enhanced while the effects on the remaining peaks are negligible. In addition, we find that the calculations without the neutrino spectral distortions cannot capture the right behavior of the matter power spectrum at small scales, where the results without the spectral distortions predict a near constant suppression at the scales $k\gtrsim 1\,h/{\rm Mpc}$, while the preciser results with the spectral distortions show that the matter power spectrum receives larger suppression at these scales. However, the discussion here refers only to the linear matter power spectrum, and we expect that a comprehensive study of the effects of neutrino spectral distortions on the LSS in the nonlinear regime can be done in the future with the N-body simulations.

### v2

@@ 5.1.2 Effects on the Neutrino Boltzmann Hierarchies
To obtain accurate CMB spectral power and matter power spectra, we must solve the coupled Boltzmann-Einstein equations for perturbations in the metric, photons, baryons, cold dark matter, and massive neutrinos. Here, we briefly review how neutrino asymmetries and spectral distortions affect the Boltzmann-Einstein equations. Following Ref. Ma:1995ey , the perturbed distribution functions for the three mass eigenstate neutrinos can be written as
@@
In this work, we adopt the synchronous gauge, where the line element is written as
$ds^{2}=a^{2}(\tau)\left\{-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\right\},$ (92)
with the metric perturbation $h_{ij}$ decomposed as $h_{ij}=h\delta_{ij}+h_{ij}^{\parallel}+h_{ij}^{\perp}+h_{ij}^{T}$, where $h\equiv h_{ii}$ is the trace of $h_{ij}$, and $h_{ij}^{\parallel}$, $h_{ij}^{\perp}$, and $h_{ij}^{T}$ are the traceless components, satisfying
@@
By definition, $h_{ij}^{\parallel}$ and $h_{ij}^{\perp}$ can be written in terms of a scalar field $\mu$ and a divergenceless vector $\vec{A}$ as
$h_{ij}^{\parallel}$ $=\left(\partial_{i}\partial_{j}-\frac{1}{3}\delta_{ij}\nabla^{2}\right)\mu,$ (94)
@@
Thus, the scalar modes of the metric perturbations are characterized by the two scalar fields $h$ and $\mu$, while the vector and tensor modes are characterized by $\vec{A}$ and $h_{ij}^{T}$, respectively. In this work, we focus solely on the scalar modes of $h_{ij}$, which can be Fourier transformed as
@@
where $h(\vec{k},\tau)$ and $\eta(\vec{k},\tau)$ represent the trace and traceless parts of the scalar modes in Fourier space, related to the scalar fields $h$ and $\mu$ in real space.
Since neutrinos are collisionless after decoupling, the evolution of $\Psi_{i}$ follows the first-order collisionless Boltzmann equation in Fourier space:
@@
where $\epsilon\equiv aE=\sqrt{q^{2}+a^{2}m^{2}_{\nu_{i}}}$ is the comoving energy. Since Eq. (96) is independent of the azimuthal angle, we can expand $\Psi_{i}$ in a Legendre series:
@@
Substituting Eq. (97) into Eq. (96) and using the orthonormality of Legendre polynomials, we obtain the Boltzmann hierarchies for $\Psi_{i,l}$:
@@
The effect of neutrino asymmetries and spectral distortions manifests through the neutrino density distribution term $\frac{d\ln\varrho_{i}^{0}}{d\ln q}$. Due to primordial neutrino asymmetries, the density matrices for neutrinos and antineutrinos of different flavors differ, leading to distinct Boltzmann hierarchies for the perturbations $\Psi_{i}$ and $\bar{\Psi}_{i}$, which must be treated separately.
With the moments $\Psi_{i,l}$, perturbations in physical quantities such as energy density, pressure, energy flux, and shear stress can be computed straightforwardly by integrating different moments. Specifically, the fractional density perturbation for neutrinos of family $i$ is given by
@@
where $\rho_{i,0}(t)$ represents the zeroth-order homogeneous background energy density for neutrinos. The effects of neutrino masses and asymmetries across different length scales are reflected in the evolution of the Fourier modes $\delta_{i,k}$, which we obtain via the CLASS code.
@@ 5.1.3 Effects on the BAO
Following Ref. Ma:1995ey , the evolution of photon perturbations is governed by the Boltzmann hierarchy:
$\dot{\delta}_{\gamma}$ $=$ $-\frac{4}{3}\theta_{\gamma}-\frac{2}{3}\dot{h},$
@@
$\dot{F}_{\gamma\,l}$ $=$ $\frac{k}{2l+1}\left[lF_{\gamma\,(l-1)}-(l+1)F_{\gamma\,(l+1)}\right]-an_{e}\sigma_{T}F_{\gamma\,l},\quad l\geq 3,$
@@
Here, $\delta_{\gamma}$ is the photon density fluctuation, $\theta_{\gamma}$ is the divergence of photon velocity, and $F_{\gamma\,l}$ represents the photon density perturbation. Additionally, $n_{e}$ is the electron number density, and $\sigma_{T}$ is the Thomson scattering cross section.
The Boltzmann equations for baryons are
$\dot{\delta}_{b}$ $=$ $-\theta_{b}-\frac{1}{2}\dot{h},$
$\dot{\theta}_{b}$ $=$ $-\frac{\dot{a}}{a}\theta_{b}+c_{s}^{2}k^{2}\delta_{b}+\frac{4\bar{\rho}_{\gamma}}{3\bar{\rho}_{b}}an_{e}\sigma_{T}(\theta_{\gamma}-\theta_{b}),$ (101)
where $\delta_{b}$ and $\theta_{b}$ are the baryon density fluctuation and the divergence of baryon fluid velocity, respectively, and $c_{s}^{2}$ is the square of the baryon sound speed. Due to electromagnetic interactions between photons and electrons, the evolution of photons and baryons differs before and after recombination. Prior to recombination, photons and baryons are tightly coupled and can be treated as a single photon-baryon fluid, known as the "tight-coupling approximation." During this "tight-coupling epoch," the perturbations of baryons and photons undergo harmonic motion with slowly decaying amplitude (Peebles:1970ag, ; Sunyaev:1970bma, ). After recombination, photons decouple from baryons, allowing the baryon acoustic oscillations (BAO) to freeze, and baryon perturbations grow as $\propto a$. The BAO feature is then imprinted on the gravitational potential and matter fluctuations, influencing the galaxy power spectrum, which depends on the sound horizon at decoupling and the angular diameter distance to the last-scattering surface.
Primordial neutrino asymmetries modify the matter-radiation equality via their contribution to $N_{\rm eff}$, altering the BAO properties. Moreover, neutrino asymmetries affect the evolution of neutrino perturbations, which in turn influence photon and baryon evolution through metric perturbations encoded in $h$ and $\eta$, despite the collisionless nature of neutrinos. As a result, primordial neutrino asymmetries significantly impact both the CMB angular power spectrum and the BAO in the matter power spectrum, as we will explore later.
@@ 5.2 Results of the CMB Anisotropies and the Matter Power Spectrum
Figure 7: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ in the SM for different total neutrino masses, compared to massless neutrinos. Right panel: The linear matter power spectrum $P(k)$ in the SM for different total neutrino masses, compared to massless neutrinos.
This subsection presents the numerical results for the CMB and LSS obtained using the CLASS code. Since both observables are influenced by neutrino masses and asymmetries, we first analyze the effect of neutrino masses on the CMB and LSS within the SM.
@@
In Fig. 7, we show the ratio of the CMB angular power spectrum, $C_{l}^{\rm TT}/C_{l}^{\rm TT}(\sum m_{\nu}=0)$, and the matter power spectrum, $P_{k}/P_{k}(\sum m_{\nu}=0)$, for several choices of total neutrino mass. The left and right panels depict the results for the CMB and matter power spectra, respectively. This analysis assumes the normal hierarchy for neutrino masses, where $\mu_{\nu_{1}}$ is adjusted for each total mass, while $\mu_{\nu_{2}}$ and $\mu_{\nu_{3}}$ are derived from the fixed mass differences between neutrinos. Focusing on the left panel, the ratio of $C_{l}^{\rm TT}$ between massive and massless neutrinos can be categorized into three stages. At large angular scales (lower multipoles, $\ell\lesssim 20$), massive neutrinos cause a small suppression due to the late-time integrated Sachs-Wolfe effect. At intermediate scales ($20\lesssim\ell\lesssim 500$), dominated by acoustic oscillations in the baryon-photon fluid, the first peak of $C_{l}^{\rm TT}$ is influenced by the modified gravitational potential evolution, resulting from the interaction between massive neutrinos and metric perturbations. Notably, the height of the first peak, due to the early integrated Sachs-Wolfe effect, is enhanced for larger neutrino masses compared to massless neutrinos. At smaller scales ($\ell\gtrsim 500$), the acoustic peaks are similarly affected, with their height increasing for larger neutrino masses.
@@
The matter power spectrum with massive neutrinos can also be divided into three stages. At large scales ($k\lesssim 10^{-3}\,h/{\rm Mpc}$), neutrinos remain outside the horizon, and the suppression of $P(k)$ results from the shift in the matter-radiation equality, caused by the transition of neutrinos to non-relativistic speeds. At intermediate scales ($10^{-3}\,h/{\rm Mpc}\lesssim k\lesssim 1\,h/{\rm Mpc}$), neutrinos gradually enter the horizon, and their free-streaming produces additional suppression in the matter power spectrum, with $P_{k}/P_{k}(\sum m_{\nu}=0)$ decreasing as $k$ increases. Finally, at small scales ($k\gtrsim 1\,h/{\rm Mpc}$), most neutrinos freely stream, leading to an almost constant suppression of the power spectrum. The larger the total neutrino mass, the greater the suppression. It is worth noting that the matter power spectrum is more sensitive to neutrino masses than the CMB, as shown in the plots. Thus, LSS observations are more effective in constraining neutrino masses, a point that will be further discussed in the next section.
@@
Figure 8: Left panel: The CMB angular power spectrum $C_{l}^{\rm TT}$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the result for $\xi_{\nu}=0$. Right panel: The linear matter power spectrum $P(k)$ for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the result for $\xi_{\nu}=0$. The total neutrino mass is fixed at $\sum m_{\nu}=0.1\,\rm{eV}$ in both panels.
@@
Figure 9: The sound horizon at the baryon drag epoch, $r_{d}$, as a function of the primordial neutrino asymmetry $\xi_{\nu}$.
@@
We now present the results for non-zero $\xi_{\nu}$. To isolate the effects of neutrino asymmetry from the total neutrino mass, the left panel of Fig. 8 shows the ratio of the CMB angular power spectrum for $\xi_{\nu}\neq 0$ to that for $\xi_{\nu}=0$, with $\sum m_{\nu_{i}}=0.1\,\text{eV}$. The corresponding ratio for the matter power spectrum is shown in the right panel. As seen in the left panel, the effects of neutrino asymmetry on the CMB angular power spectrum are primarily observed at small scales ($\ell\gtrsim 400$). For $\ell\lesssim 400$, $C_{l}^{\rm TT}$ remains nearly identical for both $\xi_{\nu}\neq 0$ and $\xi_{\nu}=0$. However, as $\ell$ increases, deviations due to $\xi_{\nu}$ become apparent. These deviations depend on both the magnitude and sign of $\xi_{\nu}$: positive $\xi_{\nu}$ enhances the small-scale power, while negative $\xi_{\nu}$ leads to suppression. This behavior is primarily driven by changes in helium abundance, which affect the CMB spectrum through diffusion damping. In addition to the overall enhancement or suppression, the acoustic peaks are also influenced by $\xi_{\nu}$, as larger values modify the oscillations at the last scattering surface. The amplitude of these modifications increases with the absolute value of $\xi_{\nu}$, with negative $\xi_{\nu}$ producing more pronounced effects.
For the linear matter power spectrum, shown in the right panel of Fig. 8, we observe that at large scales, the results for non-zero $\xi_{\nu}$ are identical to those for $\xi_{\nu}=0$. At smaller scales, however, the free-streaming of neutrinos is enhanced with increasing $\xi_{\nu}$, resulting in a slightly stronger suppression of the matter power spectrum, while this suppression is largely insensitive to the sign of $\xi_{\nu}$. A more notable effect of non-zero $\xi_{\nu}$ is the oscillations in the matter power spectrum on scales $10^{-1}\,h/{\rm Mpc}\gtrsim k\gtrsim 10^{0}\,h/{\rm Mpc}$, which reflect the influence of the BAO. Similar to the acoustic oscillations in the CMB angular power spectrum, the amplitude of BAO in the matter power spectrum increases with the absolute value of $\xi_{\nu}$, with negative $\xi_{\nu}$ leading to a larger effect than positive $\xi_{\nu}$.
@@
In addition to BAO features in the matter power spectrum, Fig. 9 shows the sound horizon at the baryon drag epoch, $r_{d}$, as a function of the primordial neutrino asymmetry $\xi_{\nu}$. Due to the additional contribution to $N_{\rm eff}$ for non-zero $\xi_{\nu}$, as per Eq. 63, larger $\left|\xi_{\nu}\right|$ corresponds to a smaller $r_{d}$, resulting in significant imprints on the matter power spectrum. These results imply that BAO experiments can provide stringent constraints on the primordial neutrino asymmetry, as will be discussed in Section 6, a consideration often overlooked in previous studies. Note that these conclusions are drawn with a fixed total neutrino mass of $\sum m_{\nu_{i}}=0.1\,\text{eV}$; however, they remain valid for different total neutrino masses, with only slight variations in the ratios $C_{l}^{\rm TT}/C_{l}^{\rm TT,\,\xi_{\nu}=0}$ and $P(k)/P^{\xi_{\nu}=0}(k)$ at intermediate scales due to the non-relativistic transition of massive neutrinos.
@@
Figure 10: The CMB angular power spectrum $C_{l}^{\rm TT}$ (left panel) and the linear matter power spectrum $P(k)$ (right panel) for different neutrino degeneracy parameters $\xi_{\nu}$, compared to the corresponding results with $\xi_{\nu}=0$. The total neutrino mass is fixed to $\sum m_{\nu}=0.1\,\text{eV}$ for both panels. The solid curves are identical to those in Fig. 5, representing the results including spectral distortions of the neutrinos and antineutrinos, while the dashed curves represent results without spectral distortions.
@@
Finally, we briefly discuss the effects of spectral distortions of neutrinos and antineutrinos on the CMB and LSS. As in Fig. 8, we present in the left and right panels of Fig. 10 the CMB angular power spectrum $C_{l}^{\rm TT}$ and the linear matter power spectrum $P(k)$ for non-zero $\xi_{\nu}$, compared to the corresponding results with $\xi_{\nu}=0$. The solid curves correspond to results that include spectral distortions, as shown in Fig. 8, while the dashed curves correspond to results without the neutrino spectral distortions. Comparing $C_{l}^{\rm TT}$ with and without spectral distortions, we see that the primary effect of including these distortions is a slight suppression of the first three acoustic peaks, with negligible impact on the smaller-scale peaks. For the linear matter power spectrum, the effect on the BAO is similar to that on the CMB spectrum, with the first two acoustic peaks slightly enhanced while the effects on subsequent peaks remain minimal. Furthermore, we find that calculations that exclude spectral distortions fail to accurately capture the behavior of the matter power spectrum at small scales. Without distortions, the results predict a near-constant suppression at $k\gtrsim 1\,h/{\rm Mpc}$, whereas more precise calculations including distortions show greater suppression at these scales. However, this analysis is limited to the linear matter power spectrum. A comprehensive study of the impact of neutrino spectral distortions on LSS in the nonlinear regime would require future N-body simulations, which is beyond the scope of this paper and left for future work.